# On Softmax Direct Preference Optimization for Recommendation

### v1

@@ 4.2 Study on S-DPO (RQ2)
To investigate the effect of explicit ranking optimization and multiple negative samples of S-DPO, we compare it with Supervised Fine-Tuning(SFT), and a variant of S-DPO with only a single negative sample, downgrading to pairwise DPO loss. The experimental results are reported in Figure 2(a). We can observe that DPO can achieve an overall better performance compared to SFT, which underscores the effectiveness of instilling ranking relationships into existing LM-based recommenders. With a more effective ranking gradient provided by multiple negative samples, S-DPO can further boost performance and achieve the best among all baseline methods and variants.
@@
Study on number of negative samples.
Benefiting from the utilization of multiple negative pairs in preference data, our S-DPO offers two empirically appealing properties compared to DPO: 1) S-DPO has more effective gradients facilitating the optimization; 2) S-DPO mitigates the data likelihood decline issue of DPO. Figure 2(b) provides the comparison of validation loss between S-DPO and DPO, illustrating that the loss of S-DPO decreases faster and more significantly. This observation demonstrates that multiple negative pairs provide larger and more meaningful gradients for model optimization, which is attributed to the inherent ability of S-DPO to mine negative samples [37] (cf. Section 3.1). More results of loss analysis can be found in Appendix C.
On the other hand, it is widely accepted that the optimization of DPO suffers from data likelihood decline issues [52, 53], which implies that the log-likelihood of the preferred completions is reduced below the original log-likelihood from the reference model, hindering the performance of DPO. We, therefore, study the behavior of S-DPO and surprisingly find that it has the potential property of mitigating data likelihood decline issues. As illustrated in Figure 2(c), S-DPO exhibits continually increasing rewards of preferred items, while DPO struggles to increase the reward of preferred items.
To further verify the superiority of the multiple negative samples of S-DPO compared with DPO, we conduct experiments to explore the potential of the number of negative samples, with the results depicted in Figure 3(a). It can be observed that utilizing multiple negative samples allows the model to achieve better performance than with a single one. Furthermore, as the number of negative samples increases, the model’s performance exhibits continual improvements. We attribute this success of S-DPO to more effective ranking gradients provided by multiple negatives which can be connected to the superior performance of contrastive loss in self-supervised recommendations [37, 38, 54]. More results of preferred item rewards can be found in Appendix C.
Study on values of $\beta$.
In S-DPO, $\beta$ is a hyperparameter controlling the deviation of LM from the base reference policy [33]. Typically, a smaller value of $\beta$ implies that the language model is more heavily influenced by the preference signals and vice versa. In this section, we explore the effect of $\beta$ on S-DPO. As indicated in Figure 3(b) through 3(c), a higher $\beta$ can achieve overall better performance in our task, while a lower $\beta$ may overwhelm the model’s learned knowledge from the supervised fine-tuning stage, as evidenced by both low valid ratio and hit ratio. On the other hand, an excessively large $\beta$ prevents the model from effectively learning ranking relationships, leading to suboptimal performance. In all our main experiments and studies, we set $\beta$ as 1 to achieve a balance between ranking signals and inherent knowledge of language models.
@@ 5.1 LM for Recommendation
Recent advancements in recommendation systems have increasingly incorporated Language Models (LMs) due to their extensive knowledge and robust reasoning abilities. This integration occurs primarily in two forms: LM-enhanced recommenders and LM-based recommenders. LM-enhanced recommenders utilize LM embedding as semantic representations to provide contrastive signals [55, 56, 57] or utilize LM as advanced feature extractors improving the representation of user and item features [58, 59, 60]. However, these systems still rely on traditional recommenders for the final recommendation task, which leaves the reasoning ability of LM largely untouched.
On the other hand, LM-based recommenders directly employ LMs for making recommendations. Early works leverage LMs’ in-context learning capabilities for zero-shot or few-shot recommendations, demonstrating significant potential [9, 10, 11, 14]. However, untuned LM-based recommenders are limited by inadequate instruction-following capabilities and a lack of domain-specific knowledge. To bridge this gap, recent efforts in this category include supervised fine-tuning of LMs on the historical interactions to enhance their performance in recommendation tasks [13, 16, 22, 23, 26]. More recently, researchers have discovered that exploring item representation methods in the finetuning phase may further boost LM’s ability for recommendation [29]. This branch of works includes integrating collaborative signals [17, 18, 24, 61], adjusting numeric representations [21, 62, 63] or introducing additional item tokens [15, 25, 27, 28].
@@ 5.2 Preference Alignment of Language Models
Reinforcement Learning from Human Feedback (RLHF) [30, 31, 32] is a prevalent method of LMs to learn from human preferences. The RLHF pipeline comprises reward model learning and reinforcement learning (RL) optimization, the latter of which suffers instability and inefficiency. Direct Preference Optimization (DPO) [33] bypasses the brittle RL phase via a particular reward model parameterization and is thus simpler to implement while still keeping the performance of RLHF.
DPO proves to be effective in many scopes, like NLP [33, 64] and multimodal LMs [34, 65, 66, 67]. Besides, several variants have been proposed for further improvement of DPO. $\Psi$PO [68] is a generalization of DPO loss and its representative IPO can better overcome the problem of overfitting. ODPO [35] treats preference pairs differently by stipulating that the likelihood gap of two responses should be greater than a corresponding offset value. Other variants like KTO [69], $f$-DPO [70], RSO [71], etc. also enhance DPO in various aspects. Despite these contributions, the possibilities for leveraging and further adapting DPO for recommendation are still largely unexplored. Moreover, there are few studies that discuss extending DPO to handle multi-negative scenarios.
@@ 6 Limitation
Despite effectiveness, there are several limitations not addressed in this paper. On the one hand, the number of negative samples is capped at 15 in our experiments. The potential of multiple negative samples hasn’t been fully explored due to the limited time and computation resources. On the other hand, increasing the number of negative examples inevitably results in higher training costs, a phenomenon that becomes more pronounced as the number of negative examples grows in the context of language models. Finally, despite the empirical success and theoretical connection to softmax loss in recommendation tasks, the profound understanding of softmax ranking loss on LM needs to be further explored.
@@ 7 Conclusion
In this work, we devised a principled Softmax-DPO (S-DPO) loss specially tailored for LM-based recommenders, utilizing multiple negatives in preference data to explicitly instill ranking information into LM. Empirically, S-DPO surpasses all baseline models including traditional and LM-based methods on three datasets in sequential recommendation tasks while successfully mitigating the data likelihood decline issue of DPO. Grounded by theoretical proof, we bridge S-DPO with the softmax loss in self-supervised recommendations, underscoring the ranking performance of S-DPO and highlighting the critical roles of multiple negatives. Also, we theoretically find that S-DPO has an inherent ability to mine hard negatives which provide larger and more effective gradients to model optimization, assuring its exceptional capabilities in recommendation tasks. We believe that our S-DPO, as a generalization of DPO, provides valuable insights for future LM-based recommenders and has the potential to benefit research fields other than recommender systems [^2] Thc extension to broader impact will be detailed in Appendix D.
@@ A.1 Deriving Preference Distribution
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in Per(ind(\mathcal{I}_{d}))}\prod_{j=1}^{K-1}\frac{{\rm exp}\left(r(x_{u},e_{\tau^{\prime}(j)})\right)}{\sum_{l=j}^{K-1}{\rm exp}(r(x_{u},e_{\tau^{\prime}(l)}))}$
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in Per(ind(\mathcal{I}_{d}))}p^{*}(\tau^{\prime}|x_{u},\mathcal{E}_{d})$
@@
wherein $ind(\mathcal{E}_{d})$ denotes the indices of titles in $\mathcal{E}_{d}$ and $Per(ind(\mathcal{E}_{d}))$ denotes the set of permutations of index set $ind(\mathcal{E}_{d})$. The third equation is because a permutation of $\{1,2\cdots,K\}$ starting with $p$ can be divided into the prefix $p$ and a subsequent permutation of the rest indices, which is exactly $Per(ind(\mathcal{E}_{d}))$.
@@ A.3 Deriving the Gradient of S-DPO Loss
Using properties of sigmoid function that $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))=\sigma(x)\sigma(-x)$ and thus $\left(({\rm log}\,\sigma(x)\right)^{\prime}=\frac{1}{\sigma(x)}\times\sigma(x)\sigma(-x)=\sigma(-x)$, we have:
@@ B.3 Implementation Details
We implement all approaches with Python 3.9.7, PyTorch 2.2.2, and transformers 4.39.0 on 4 NVIDIA A100s. We select Llama2-7B [31] as the LM backbone for S-DPO. Following [18], we randomly select prompts from several prompt formats during training and evaluation to ensure flexibility and generality. For optimization of all the traditional methods, the Adam optimizer is employed with a learning rate adjusted to 0.001, and a batch size configured at 256. All models undergo L2 regularization, with coefficients experimentally determined from [1e-3, 1e-4, 1e-5, 1e-6, 1e-7]. In all experiments involving large language models, we train each method for a maximum of 5 epochs using a batch size of 128 and select the checkpoint with the lowest loss on the validation set as the final checkpoint. A warm-up strategy is applied to the learning rate, starting at 5% of its maximum value, and gradually adjusting it through a cosine scheduler throughout the training process. For S-DPO and all of its ablation studies, we further conduct preference training for further 3 epochs with a batch size of 128 and a learning rate of 1e-5. Setting the value of $\beta$ as 1, we search the number of negative samples in [3,5] for the main results. The effects of both factors are further explored in 4.2.
@@ B.4 Evaluation Metrics
Appendix C Study of Negative Samples Number
(a) Study of valid loss on MovieLens.
(b) Study of valid loss on Goodreads.
Figure 4: Study on validation loss between S-DPO and DPO on MovieLens and Goodreads.
Figure 4(a) through figure 4(b) provide the comparison of validation loss between S-DPO and DPO on MovieLens and Goodreads, illustrating that the loss of S-DPO decreases faster and more significantly. This observation is consistent with Figure 2(b).
On the other hand, figure 5(a) through figure 5(b) showcase that S-DPO exhibits continually increasing rewards of preferred items, while DPO struggles to increase the reward of preferred items on both MovieLens and Goodreads, which is aligned with Figure 2(c).
(a) Study of preferred item rewards on MovieLens.
(b) Study of preferred item rewards on Goodreads.
Figure 5: Study on validation loss between S-DPO and DPO on MovieLens and Goodreads.

### v3

@@ 4.2 Study on S-DPO
To investigate the effect of explicit ranking optimization and multiple negative samples of S-DPO, we compare it with the vanilla supervised fine-tuned model (w/o ranking), and a variant of S-DPO with only a single negative sample (w/o multi-neg), downgrading to pairwise DPO loss. The experimental results are reported in Figure 2(a). We can observe that DPO can achieve an overall better performance compared to SFT, which underscores the effectiveness of instilling ranking relationships into existing LM-based recommenders. With a more effective ranking gradient provided by multiple negative samples, S-DPO can further boost performance and achieve the best among all baseline methods and variants.
Study on the number of negative samples (RQ2).
Benefiting from the utilization of multiple negative pairs in preference data, our S-DPO offers two empirically appealing properties compared to DPO: 1) S-DPO has more effective gradients facilitating the optimization; 2) S-DPO provides a better boost for rewards of preferred items compared to DPO. Figure 2(b) provides the comparison of validation loss between S-DPO and DPO, illustrating that the loss of S-DPO decreases faster and more significantly. This observation demonstrates that multiple negative pairs provide larger and more meaningful gradients for model optimization, which is attributed to the inherent benefit of S-DPO to mine negative samples [38] (cf. Section 3.1).
On the other hand, we study the behavior of S-DPO which is illustrated in Figure 2(c). We surprisingly find that S-DPO exhibits continually increasing rewards of preferred items that are more significant and stable than DPO, which shows better effectiveness in distinguishing preferred items and a potential of mitigating data likelihood decline issues[52, 53].
To further verify the superiority of the multiple negative samples of S-DPO compared with DPO, we select the number of negative samples from {1, 3, 5, 8, 10, 15} to conduct experiments to explore the potential of the number of negative samples, with the results depicted in Figure 3(a). It can be observed that utilizing multiple negative samples allows the model to achieve better performance than with a single one. Furthermore, as the number of negative samples increases, the model’s performance exhibits continual improvements. We attribute this success of S-DPO to more effective ranking gradients provided by multiple negatives which can be connected to the superior performance of contrastive loss in self-supervised recommendations [38, 39, 54].
To validate the superiority of S-DPO over the DPO variant with multi-negatives, we conduct effectiveness and efficiency comparisons. Table 2 demonstrates that introducing more negative samples benefits both DPO and S-DPO, and S-DPO achieves comparable performance with fewer training steps. We further analyze how S-DPO outperforms DPO in terms of computational efficiency. While the complexity of DPO for $K$ negative samples is $\Theta(2KC_{\mathcal{M}}S_{t})$, where $C_{\mathcal{M}}+1$ denotes the base LLM’s computational complexity and $S_{t}$ represents the size of the training data, S-DPO’s complexity for the same number of negatives is reduced to $\Theta((K+1)(C_{\mathcal{M}}+1)S_{t})$. This efficiency gain can be expressed by scaling the complexity of DPO by the factor $\frac{1}{2}+\frac{1}{2K}+\frac{1}{2C_{\mathcal{M}}}+\frac{1}{2KC_{\mathcal{M}}}$, highlighting that S-DPO offers significant advantages, especially when working with a larger number of negative samples.
Table 2: Effectiveness comparison between DPO with single negative, a variant of DPO with multiple negatives and S-DPO with the same number of negatives (we set $K$ as 3 to get the performance in this table).
. Datasets LastFM MovieLens Goodreads Complexity Measure HitRatio@1 ValidRatio HitRatio@1 ValidRatio HitRatio@1 ValidRatio DPO-$1$negative 0.6342 0.9972 0.4947 0.9684 0.6381 0.9900 $\Theta(2C_{\mathcal{M}}S_{t})$ DPO-$K$negative 0.6413 0.9964 0.4947 0.9474 0.6628 0.9900 $\Theta(2KC_{\mathcal{M}}S_{t})$ S-DPO-$K$negative 0.6477 0.9980 0.5263 0.9895 0.6661 0.9950 $\Theta((K+1)(C_{\mathcal{M}}+1)S_{t})$
@@
Study on values of $\beta$ (RQ3).
In S-DPO, $\beta$ is a hyperparameter controlling the deviation of LM from the base reference policy [34]. Typically, a smaller value of $\beta$ implies that the language model is more heavily influenced by the preference signals and vice versa. In this section, we select the value of $\beta$ from {0.1, 0.5, 1, 3, 5} to explore the effect of $\beta$ on S-DPO. As indicated in Figure 3(b) through 3(c), a higher $\beta$ can achieve overall better performance in our task, while a lower $\beta$ may overwhelm the model’s learned knowledge from the supervised fine-tuning stage, as evidenced by both low valid ratio and hit ratio. On the other hand, an excessively large $\beta$ prevents the model from effectively learning ranking relationships, leading to suboptimal performance. In all our main experiments and studies, we set $\beta$ as 1 to achieve a balance between ranking signals and inherent knowledge of language models.
@@ 5.1 LM for Recommendation
Recent advancements in recommendation systems have increasingly incorporated Language Models (LMs) due to their extensive knowledge and robust reasoning abilities. This integration occurs primarily in two forms: LM-enhanced recommenders and LM-based recommenders. LM-enhanced recommenders utilize LM embedding as semantic representations to provide contrastive signals [55, 56, 57, 58] or utilize LM as advanced feature extractors improving the representation of user and item features [59, 60, 61]. However, these systems still rely on traditional recommenders for the final recommendation task, which leaves the reasoning ability of LM largely untouched.
On the other hand, LM-based recommenders directly employ LMs for making recommendations. Early works leverage LMs’ in-context learning capabilities for zero-shot or few-shot recommendations, demonstrating significant potential [10, 11, 12, 15]. However, untuned LM-based recommenders are limited by inadequate instruction-following capabilities and a lack of domain-specific knowledge. To bridge this gap, recent efforts in this category include supervised fine-tuning of LMs on the historical interactions to enhance their performance in recommendation tasks [14, 17, 23, 24, 27]. More recently, researchers have discovered that exploring item representation methods in the finetuning phase may further boost LM’s ability for recommendation [30]. This branch of works includes integrating collaborative signals [18, 19, 25, 62, 63, 64], adjusting numeric representations [22, 65, 66] or introducing additional item tokens [16, 26, 28, 29].
@@ 5.2 Preference Alignment of Language Models
Reinforcement Learning from Human Feedback (RLHF) [31, 32, 33] is a prevalent method of LMs to learn from human preferences. The RLHF pipeline comprises reward model training and reinforcement learning (RL) optimization, the latter of which suffers instability and inefficiency. Direct Preference Optimization (DPO) [34] bypasses the brittle RL phase via a particular reward model parameterization and is thus simpler to implement while still keeping the performance of RLHF.
DPO proves to be effective in many scopes, like NLP [34, 67] and multimodal LMs [35, 68, 69, 70]. Besides, several variants have been proposed for further improvement of DPO. $\Psi$PO [71] is a generalization of DPO loss and its representative IPO can better overcome the problem of overfitting. ODPO [36] treats preference pairs differently by stipulating that the likelihood gap of two responses should be greater than a corresponding offset value. KTO [72] utilizes prospect theory for preference alignment tasks. Other variants including GPO [73], $f$-DPO [74], RSO [75] also enhance or expand DPO in various aspects. Despite these contributions, the possibilities for leveraging and further adapting DPO for recommendation are still largely unexplored and few studies discuss extending DPO to handle multi-negative scenarios.
@@ 6 Limitation
Despite effectiveness, there are several limitations not addressed in this paper. On the one hand, the number of negative samples is capped at 15 in our experiments. The potential of multiple negative samples hasn’t been fully explored due to the limited time and computation resources. On the other hand, increasing the number of negative examples inevitably results in higher training costs, a phenomenon that becomes more pronounced as the number of negative examples grows in the context of language models.
@@ 7 Conclusion
In this work, we devised a principled Softmax-DPO (S-DPO) loss specially tailored for LM-based recommenders, utilizing multiple negatives in preference data to explicitly instill ranking information into LM. Empirically, S-DPO surpasses all baseline models including traditional and LM-based methods on three datasets in sequential recommendation tasks while successfully providing better rewards for preferred items compared to DPO. Grounded by theoretical proof, we bridge S-DPO with the softmax loss in self-supervised recommendations, underscoring the ranking performance of S-DPO and highlighting the critical roles of multiple negatives. Also, we theoretically find that S-DPO has an inherent benefit to mine hard negatives which provide larger and more effective gradients to model optimization, assuring its exceptional capabilities in recommendation tasks. We believe that S-DPO, as a generalization of DPO, provides valuable insights for future LM-based recommenders and has the potential to benefit research fields other than recommender systems [^2] Thc extension to broader impact will be detailed in Appendix D.
Acknowledgments and Disclosure of Funding
This research is supported by the National Science and Technology Major Project (2023ZD0121102), the NExT Research Center, National Natural Science Foundation of China (92270114). This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ A.1 Deriving Preference Distribution
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in\mathrm{Per}(\mathrm{ind}(\mathcal{I}_{d}))}\prod_{j=1}^{K-1}\frac{{\rm exp}\left(r(x_{u},e_{\tau^{\prime}(j)})\right)}{\sum_{l=j}^{K-1}{\rm exp}(r(x_{u},e_{\tau^{\prime}(l)}))}$
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in\mathrm{Per}(\mathrm{ind}(\mathcal{I}_{d}))}p^{*}(\tau^{\prime}|x_{u},\mathcal{E}_{d})$
@@
wherein $\mathrm{ind}(\mathcal{E}_{d})$ denotes the indices of titles in $\mathcal{E}_{d}$ and $\mathrm{Per}(\mathrm{ind}(\mathcal{E}_{d}))$ denotes the set of permutations of index set $\mathrm{ind}(\mathcal{E}_{d})$. The third equation is because a permutation of $\{1,2\cdots,K\}$ starting with $p$ can be divided into the prefix $p$ and a subsequent permutation of the rest indices, which is exactly $\mathrm{Per}(\mathrm{ind}(\mathcal{E}_{d}))$.
@@ A.3 Deriving the Gradient of S-DPO Loss
Using properties of the sigmoid function that $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))=\sigma(x)\sigma(-x)$ and thus $\left(({\rm log}\,\sigma(x)\right)^{\prime}=\frac{1}{\sigma(x)}\times\sigma(x)\sigma(-x)=\sigma(-x)$, we have:
@@ B.3 Implementation Details
We implement all approaches with Python 3.9.7, PyTorch 2.2.2, and transformers 4.38.2 on 4 NVIDIA A100s. We select Llama2-7B [32] as the LM backbone for S-DPO. Following [19], we randomly select prompts from several prompt formats during training and evaluation to ensure flexibility and generality. For optimization of all the traditional methods, the Adam optimizer is employed with a learning rate adjusted to 0.001, and a batch size configured at 256. All models undergo L2 regularization, with coefficients experimentally determined from [1e-3, 1e-4, 1e-5, 1e-6, 1e-7]. In all experiments involving large language models, we train each method for a maximum of 5 epochs using a batch size of 128 and select the checkpoint with the lowest loss on the validation set as the final checkpoint. A warm-up strategy is applied to the learning rate, starting at 5% of its maximum value, and gradually adjusting it through a cosine scheduler throughout the training process. For S-DPO and all of its ablation studies, we further conduct preference training for further 3 epochs with a batch size of 128 and a learning rate of 1e-5. Setting the value of $\beta$ as 1, we search the number of negative samples in [3,5] for the main results. The effects of both factors are further explored in 4.2.
@@ B.4 Evaluation Metrics
Appendix C Study on Backbone Language Models
Table 4: The performance comparison among three different backbone language models on LastFM and MovieLens.
LLaMA1-7B Mistral-7B Pythia-2.8B
HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio
LastFM Vanilla 0.0465 0.5872 0.0633 0.3648 0.0265 0.3648
Language Modeling 0.5980 0.9980 0.7828 0.9992 0.1611 0.4281
DPO 0.6084 0.9976 0.7415 0.9964 0.1896 0.4220
S-DPO (3 negatives) 0.6285 0.9976 0.7679 0.9972 0.1948 0.4689
S-DPO (8 negatives) 0.6365 0.9988 0.7820 0.9972 0.2200 0.4685
MovieLens Vanilla 0.0316 0.5158 0.0842 0.6737 0.0421 0.4421
Language Modeling 0.3895 0.9684 0.4211 0.9895 0.1053 0.5684
DPO 0.3789 0.9684 0.4421 0.9684 0.1271 0.8449
S-DPO (3 negatives) 0.4526 0.9474 0.4421 0.9895 0.1271 0.8737
S-DPO (8 negatives) 0.4526 0.9579 0.4947 0.9895 0.1474 0.8737
In order to examine whether the superiority of S-DPO loss over traditional language modeling loss can be generalized across different backbone language models, we conducted experiments using models with varying architectures and sizes, including LLAMA1-7b [77], Pythia-2.8b [78], and Mistral-7b [79]. These models were tested on two distinct datasets: LastFM and MovieLens. We evaluated three training approaches: language models fine-tuned with standard language modeling loss, and models further trained using DPO and S-DPO losses. Due to limitations in computational resources and time, we experimented with two variants of S-DPO: S-DPO (3 negatives) and S-DPO (8 negatives). As shown in Table 4, S-DPO consistently outperformed the language modeling loss, enhancing model performance while maintaining or even improving the validity of the generated outputs. Additionally, the performance continued to improve as the number of negative samples increased from 3 to 8.